OpenStudy (anonymous):

I just need someone to check if I did this right.Find the distance between (-2, 8) and (11, 2). Round to the nearest hundredth. The answer I got was 89. Did I do anything wrong?

OpenStudy (anonymous):

seems unlikely for the distance to be that large did you forget the square root?

OpenStudy (anonymous):

I'm supposed to square 89, aren't I..

OpenStudy (anonymous):

I now got 9.43...

jimthompson5910 (jim_thompson5910):

how did you get 89?

OpenStudy (anonymous):

Well.. I used the distance formula

jimthompson5910 (jim_thompson5910):

so you would have done subtract x coordinates: 11 - (-2) = 11 + 2 = 13 subtract y coordinates: 2 - 8 = -6 then square each difference: 13^2 = 169 and (-6)^2 = 36 after that you add the two squares: 169 + 36 = 205

jimthompson5910 (jim_thompson5910):

the last step is to take the square root of 205 to get sqrt(205) = 14.3178210632763 roughly that rounds to 14.32

jimthompson5910 (jim_thompson5910):

hopefully you can spot where you went wrong

OpenStudy (anonymous):

yeah i see it. I didn't change the negative 2 into a positive

jimthompson5910 (jim_thompson5910):

ah glad you caught your error

OpenStudy (anonymous):

yeah ha. I remember doing this but I guess it just slipped my mind thank you for pointing that out to me. I would've never caught it

jimthompson5910 (jim_thompson5910):

you're welcome